OpenStudy (anonymous):

For problems 11–12, find the distance between the two points. Round to the nearest tenth if necessary. 11. (8, 8), (12, 11) (1 point) 7 5 25 28 12. (–2, –6), (3, 9) (1 point) 15.8 250 3.2 20

OpenStudy (anonymous):

11. d = \[\sqrt{(12-8)^2 + (11-8)^2}\] = 5 12. d = \[\sqrt{(9+6)^2 + (3+2)^2}\] = \[\sqrt{250}\]